### v1

@@ Acknowledgements
[…]
Author contributions
[…]
Code availability
All our code is open-sourced under the MIT license and can be found at:
[…]
Data availability
The dataset used can be found at https://huggingface.co/datasets/BioDT/BioCube and its code base at https://github.com/BioDT/bfm-data.
Appendix A BioAnalyst Implementation Details
@@ A.1 Core Architectural Blueprint
Firstly, $\mathbf{X}_{t}$ is first processed by an encoder $\mathcal{E}$ module built upon Perceiver IO, which transforms the input data into a fixed-size latent representation $\mathbf{Z}_{t}\in\mathbb{R}^{N_{l}\times D_{e}}$, where $N_{l}$ is the number of latent tokens and $D_{e}$ is the embedding dimensions. The encoder can process inputs from one or more time steps (e.g., $t$ and $t-1$) to form $\mathbf{Z}_{t}$. This stage includes methods, including positional and temporal encodings, as detailed in later sub-appendices.
@@
A.2 The BioAnalyst Encoder
@@
A.2.1 Input Processing and Feature Engineering
@@
  • Spatial coordinate encoding: the normalized centroid coordinates $(x,y)$ (i.e., $x,y\in[-1,1]$)) for each patch are encoded using Fourier features. For $N_{f}$ frequency bands (which in our case were set to $N_{f}=64$ and a maximum frequency of 224), the features for each coordinate are $[\sin(s_{0}\pi x),\cos(s_{0}\pi x),\dots,\sin(s_{N_{f}-1}\pi x),\cos(s_{N_{f}-1}\pi x)]$, where $s_{k}$ are linearly spaced frequencies. These are concatenated with the original coordinates and projected to the model’s embedding dimension $D_{e}$.
@@
A.2.2 Perceiver IO Latent Transformation
@@ A.3 The BioAnalyst Backbone
A.3.1 Backbone Encoder Path
@@
A.3.2 Backbone Decoder Path
@@
Each stage then applies Swin Transformer blocks. The final decoder output, $\mathbf{Z}’_{t+1}$, matches the input $\mathbf{Z}_{t}$ in shape ($N_{l}\times D_{e}$) and represents the predicted next latent state.
A.3.3 Hierarchical Processing with Shifted Windows
@@ A.4 The BioAnalyst Decoder
A.4.1 Output Query Formulation
@@
After attending to $\mathbf{Z}^{\prime}_{t+1}$, the decoder produces output embeddings of shape $N_{q}\times D_{e}$. These are projected through task-specific layers into flat variable maps, then reshaped into ×$\mathcal{H}\times\mathcal{W}$ grids—one for each target variable, level, or species.
A.4.2 Cross-Attention with Backbone Output
@@
A.4.3 Projection and Reshaping to Final Output
@@
A.5 Positional, Temporal, and Variable-Specific Encodings
@@
A.5.1 Temporal Encoding Schemes
@@
A.5.2 Variable-Specific and Categorical Feature Embeddings
@@ B.1 Introduction
BioAnalyst is pre-trained on a part of BioCube as discussed at section 3.2, using a Batch structure throughout with statistics available at Table 5.
@@ B.2 Variable groups
For our data mixture we have used the below variable groups $V$ with their corresponding variables $v$:
  • Surface variables: t2m, msl, slt, z, u10, v10, lsm
  • Edaphic variables: swvl1, swvl2, stl1, stl2
  • Atmospheric variables: z, t, u, v, q
  • Pressure levels: 1000, 925, 850, 700, 600, 500, 400, 300, 250, 200, 150, 100, 50
  • Climate variables: smlt, tp, csfr, avg_sdswrf, avg_snswrf, avg_snlwrf, avg_tprate, avg_sdswrfcs, sd, t2m, d2m
  • Miscellaneous variables: avg_slhtf, avg_pevr
  • Vegetation variables: NDVI
  • Land variables: Land (percentage of total land area)
  • Agriculture variables: Agriculture, Arable, Cropland (percentage of land area)
  • Redlist variables: RLI (indicator of the changing state of global biodiversity)
  • Forest variables: Forest (percentage of land area)
  • Species variables: Species occurrences records
@@
More specific, climatic energy and moisture variables like temperature and humidity reflect thermal nichies where humidity gives vapour-pressure deficit a key plant and anthropogenic stressor [44]. Precipitation indicates the water balance which is the strongest global predictor after temperature [45]. Radiation on various lengths (short and long) drives photosynthesis and evapotranspiration and has showed improvement in Net Primary Productivity (NPP)-linked biodiversity models [46]. Snow variables at high latitudes and mountains control growing season length and habitat suitability.
For edaphic water and temperature status, we selected to look into soil moisture which is a direct proxy for plant water stress and root-zone dynamics and has a strong interactive effect with species richness [47]. Soil temperature provides germination cues linked with soil microbial activity while variables like potential evapotranspiration and surface latent heat flux are need for water deficit estimation [48].
@@
Finally, for species we select 2 variables to work with. The first is the Red List Index (RLI) which is an indicator of the changing state of global biodiversity and defines the conservation status of major species group and measures trends in extinction risk over time. The second is the species occurrences records for 6 major categories that is highly ranked in importance from European Union. More specific Large carnivores, farmland birds, wild pollinators, herpetofauna, invasive alien species (IAS) and Mediterranean endemics together span the core biodiversity priorities currently driving European nature policy and funding. Large carnivore management has become a political flash-point: the Nature Restoration Regulation explicitly references “conflict species” such as the wolf, and the European Parliament voted in May 2025 to downgrade the wolf from strict to general protection under the Habitats Directive [49]. Farmland birds remain the EU’s headline biodiversity indicator; the European Environment Agency reports a $40\%$ decline in the Farmland Bird Index since 1990, underscoring the need for landscape-scale restoration [50]. Wild pollinators stand at the centre of the revised EU Pollinators Initiative (“A new deal for pollinators”), which commits all Member States to continent-wide monitoring and trend reversal by 2030 [51]. IAS require surveillance and early-warning systems under Regulation (EU) 1143/2014, which obliges Member States to establish national monitoring and rapid-response mechanisms [52]. Finally, herpetofauna and Mediterranean endemics receive targeted LIFE funding and are focal taxa in Article 17 conservation-status reporting under the Habitats Directive, anchoring them in the EU’s mandatory assessment cycle. Table 3 lists in detail all the species from the above categories that are used for training BioAnalyst.
Table 3: Approximate total occurrences between 2000-2020 and species ID to scientific name mappings according to GBIF for six major species categories
@@
Invasive & Alien Species 8002952 Ambrosia artemisiifolia (common ragweed) 87 k
@@
Mediterranean Species 2435261 Lynx pardinus (Iberian lynx) 436 k
@@
2434779 Monachus monachus (Mediterranean monk seal) 137 k
@@
Table 4: ERA5 variable names, their definitions and measure units [53] [54]
chat Variable Units Description
swvl1 m3 m–3 Volumetric soil water content, layer 1 (0cm depth)
swvl2 m3 m–3 Volumetric soil water content, layer 2 (7cm depth)
@@
smlt m (water equivalent) Snow melt accumulated at surface
@@
csfr kg m–2 s–1 Convective snowfall rate
avg_sdswrf W m–2 Mean surface downwelling shortwave radiation flux
avg_snswrf W m–2 Mean surface net shortwave radiation flux
avg_snlwrf W m–2 Mean surface net longwave radiation flux
avg_tprate m s–1 Mean total precipitation rate
avg_sdswrfcs W m–2 Mean surface downwelling shortwave flux (clear sky)
@@
u m s–1 Eastward wind component (pressure levels)
v m s–1 Northward wind component (pressure levels)
u10 m s–1 10m eastward wind component
v10 m s–1 10m northward wind component
q kg kg–1 Specific humidity (pressure levels)
lsm 0/1 Land-sea mask (0=sea, 1=land)
avg_slhtf W m–2 Mean surface latent heat flux
avg_pevr kg m–2 s–1 Mean potential evaporation rate
@@ B.3 Build Batches
Every input dataset is preprocessed and harmonized based on the following principles:
@@
  • Timestamps are standardized to datetime64 objects with monthly resolution.
@@
To maintain modularity and extensibility, the pipeline separates data loading logic per modality (e.g., _load_era5, _load_csv, _load_species), with all outputs standardized into spatially and temporally aligned xarray.Dataset objects. This design supports the seamless addition of future data types such as Sentinel-2 imagery, elevation, or anthropogenic indicators without altering the core batch assembly logic.
The batching pipeline is optimized to handle large-scale datasets efficiently. ERA5 NetCDF files are processed using xarray with chunking enabled, and CSV files are filtered and indexed spatially using precomputed $0.25^{\circ}$ grid maps. This architecture allows scaling to continental datasets and long temporal ranges without excessive memory usage.
@@
All variables are converted into PyTorch tensors. Scalar geospatial variables (e.g., temperature, NDVI) are stored in tensors of shape $(2,H,W)$, where $H$ and $W$ are the grid height and width and 2 accounts for the consecutive timestamps. Pressure-level variables (e.g., geopotential, wind at altitude) are stored as $(2,C,H,W)$ where $C$ is the number of pressure levels. Species presence tensors are organized as one tensor per species with shape $(2,H,W)$. In cases where pressure-level variables are included, their corresponding pressure levels are saved in the batch metadata.
Each batch includes a metadata dictionary capturing the timestamp window, grid specification (latitude and longitude), list of included species, and pressure levels (if applicable). The final batch is a structured dictionary with modality keys representing the various variable categories (e.g., surface_variables, agriculture_variables, species_variables), each containing variable-specific tensors. The complete batch is serialized and saved to disk in PyTorch’s binary .pt format, enabling efficient loading during model training without repeating and preprocessing operation.
Table 5: Statistics of the constructed data Batches used in the pre-training of BioAnalyst
@@
Grid Sampling Resolution 0.25°
Global Grid Size 1440 (lon) $\times$ 720 (lat) = 1,036,800 cells
Europe Grid Size 280 (lon) $\times$ 160 (lat) = 44,800 cells
Latitude Bounds [32.0°, 72.0°]
Longitude Bounds [–25.0°, 45.0°]

### v2

@@ Acknowledgments
Appendix contents
Appendix A BioAnalyst Model
@@ A.1 Core Architectural Blueprint
Firstly, $\mathbf{X}_{t}$ is first processed by an encoder $\mathcal{E}$ module built upon Perceiver IO, which transforms the input data into a fixed-size latent representation $\mathbf{Z}_{t}\in\mathbb{R}^{N_{l}\times D_{e}}$, where $N_{l}$ is the number of latent tokens and $D_{e}$ is the embedding dimensions. The encoder can process inputs from one or more time steps (e.g., $t$ and t1t-1) to form $\mathbf{Z}_{t}$. This stage includes methods, including positional and temporal encodings, as detailed in later sub-appendices.
@@
A.2 Design Choices
We have chosen Swin Transformer as our backbone because biodiversity patterns can be observed at multiple spatial scales (e.g., from microclimates to continental zones), and Swin’s hierarchical architecture with patch merging/splitting appeared as an interesting and suitable choice for a backbone that would ingest data with multi-scale structuring. Additionally, its window-based attention mechanism also is stated as having linear complexity (when compared to the original ViT) in processing image/image-like inputs, making it a computationally efficient choice - a property that certainly would be desirable in the case of processing/training/testing on large inputs (e.g., input data spanning our entire planet, not just continental Europe). Additionally, we adjusted the architecture to support temporal conditioning through lead-time embeddings that can modulate processing at each stage, allowing the model to adapt representations based on the prediction time horizon, which is a capability we use for multi-month forecasting.
Integrating Perceiver IO as encoder-decoder and Swin Transformer as backbone is achieved through a simple reshape operation that reinterprets the latent sequence as a 3D spatial volume. The Perceiver IO encoder outputs a flat sequence of latent tokens, which is then reshaped into a spatial tensor with dimensions (depth, height, width), where the depth dimensions naturally aligns with the number of modality groups and the spatial dimensions correspond to the patch grid. This creates a multi-channel feature map, similarly to standard computer vision inputs, which the Swin backbone can then process through hierarchical window-based attention on the spatial dimensions while treating depth as feature channels. After Swin processing, the inverse transformation flattens the spatial volume back into a sequential format for the Perceiver IO decoder, which uses cross-attention to generate pixel-level predictions. This design combines Perceiver IO’s strength at flexible, modality-agnostic encoding with Swin’s hierarchical processing. The depth alignment with modality groups is by design: the encoder initialises separate latent parameter sets for each modality group (each containing one latent per spatial patch), which are concatenated in a structured order before Perceiver IO processing, which ensures that the subsequent, deterministic reshaping will map each modality to its own depth slice in the 3D volume.
A.3 The BioAnalyst Encoder
@@
A.3.1 Input Processing and Feature Engineering
@@
  • Spatial coordinate encoding: the normalized centroid coordinates $(x,y)$ (i.e., $x,y\in[-1,1]$)) for each patch are encoded using Fourier features. For $N_{f}$ frequency bands (which in our case were set to $N_{f}=64$ and a maximum frequency of 224), the features for each coordinate are $\sin(s_{0}\pi x),\cos(s_{0}\pi x),\dots,\sin(s_{N_{f}-1}\pi x),\cos(s_{N_{f}-1}\pi x)]$, where $s_{k}$ are linearly spaced frequencies. These are concatenated with the original coordinates and projected to the model’s embedding dimension $D_{e}$.
@@
A.3.2 Perceiver IO Latent Transformation
@@ A.4 The BioAnalyst Backbone
A.4.1 Backbone Encoder Path
@@
A.4.2 Backbone Decoder Path
@@
Each stage then applies Swin Transformer blocks. The final decoder output, $\mathbf{Z}^{\prime}_{t+1}$, matches the input $\mathbf{Z}_{t}$ in shape ($N_{l}\times D_{e}$) and represents the predicted next latent state.
A.4.3 Hierarchical Processing with Shifted Windows
@@ A.5 The BioAnalyst Decoder
A.5.1 Output Query Formulation
@@
After attending to $\mathbf{Z}^{\prime}_{t+1}$, the decoder produces output embeddings of shape $N_{q}\times D_{e}$. These are projected through task-specific layers into flat variable maps, then reshaped into $\mathcal{H}\times\mathcal{W}$ grids—one for each target variable, level, or species.
A.5.2 Cross-Attention with Backbone Output
@@
A.5.3 Projection and Reshaping to Final Output
@@
A.6 Positional, Temporal, and Variable-Specific Encodings
@@
A.6.1 Temporal Encoding Schemes
@@
A.6.2 Variable-Specific and Categorical Feature Embeddings
@@ A.7 Data Normalization
A.8 Emulation stage
Given a system state $\mathbf{X}_{t}$ at time $t$, BioAnalyst goal is to predict the next state $\mathbf{X}_{t^{\prime}}$ at time $t^{\prime}>t$. In the common single-step forecast scenario, following (bodnar2024aurora), we define a simulator function
@@ B.1 Introduction
BioAnalyst is pre-trained on a part of BioCube as discussed at section 3.2, using a Batch structure throughout with statistics available at Table 5. Our decision to use $0.25^{\circ}$ grid resolution is motivated by a substantial body of work in macroecology and conservation planning that operates at comparable or even coarser resolutions when addressing broad-scale questions (e.g. global biodiversity risk, identification of priority regions, assessment of conservation networks) (engemann2015limited). For example, recent global analyses of climate-change risks to terrestrial biodiversity and macroecological diversity patterns have used 25–50 km grid cells, arguing that such coarse units are appropriate for capturing broad climatic and biogeographic gradients relevant to policy (roll2017global). Similarly, assessments of Key Biodiversity Areas and other large-scale conservation prioritization exercises often use 25–50 km planning units. Thus, a 0.25° grid is well aligned with the spatial grain at which many national or continental assessments are currently performed (farooq2023call).
@@ B.2 Variable groups
For our data mixture we have used the 11 variable groups $V$ with their corresponding variables $v$ available on Table 2.
Table 2: Overview of variable groups used as input features for the pre-training and fine-tuning of BioAnalyst.
Variable group Variables Dimensionality
Surface variables
t2m, msl, slt, z,
u10, v10, lsm
[160, 280, 7]
Edaphic variables
swvl1, swvl2,
stl1, stl2
[160, 280, 4]
Pressure levels (atmospheric)
$1000$, $925$, $850$, $700$, $600$, $500$,
$400$, $300$, $250$, $200$, $150$, $100$, $50$
13
Atmospheric variables
z, t, u, v, q
[160, 280, 5, 13]
Climate variables
smlt, tp, csfr, avg_sdswrf,
avg_snswrf, avg_snlwrf,
avg_tprate, avg_sdswrfcs, sd,
t2m, d2m
[160, 280, 11]
Miscellaneous variables
avg_slhtf, avg_pevr
[160, 280, 2]
Vegetation variables NDVI [160, 280, 1]
Land variables Land (percentage of total land area) [160, 280, 1]
Agriculture variables
Agriculture, Arable, Cropland
(percentage of land area)
[160, 280, 3]
Redlist variables RLI (Red List Index for biodiversity state) [160, 280, 1]
Forest variables Forest (percentage of land area) [160, 280, 1]
Species variables Species occurrence records [160, 280, 28]
@@
More specific, climatic energy and moisture variables like temperature and humidity reflect thermal nichies where humidity gives vapour-pressure deficit a key plant and anthropogenic stressor (zuquim2020importance). Precipitation indicates the water balance which is the strongest global predictor after temperature (GUTIERREZHERNANDEZ2021239). Radiation on various lengths (short and long) drives photosynthesis and evapotranspiration and has showed improvement in Net Primary Productivity (NPP)-linked biodiversity models (rs12030372). Snow variables at high latitudes and mountains control growing season length and habitat suitability.
For edaphic water and temperature status, we selected to look into soil moisture which is a direct proxy for plant water stress and root-zone dynamics and has a strong interactive effect with species richness (XU2022149950). Soil temperature provides germination cues linked with soil microbial activity while variables like potential evapotranspiration and surface latent heat flux are need for water deficit estimation (hess-28-2617-2024).
@@
Finally, for species we select 2 variables to work with. The first is the Red List Index (RLI) which is an indicator of the changing state of global biodiversity and defines the conservation status of major species group and measures trends in extinction risk over time. The second is the species occurrences records for 6 major categories that is highly ranked in importance from European Union. More specific large carnivores, farmland birds, wild pollinators, herpetofauna, invasive alien species (IAS) and Mediterranean endemics together span the core biodiversity priorities currently driving European nature policy and funding. Large carnivore management has become a political flash-point: the Nature Restoration Regulation explicitly references “conflict species” such as the wolf, and the European Parliament voted in May 2025 to downgrade the wolf from strict to general protection under the Habitats Directive (EU_NRL_Proposal_2022). Farmland birds remain the EU’s headline biodiversity indicator; the European Environment Agency reports a $40\%$ decline in the Farmland Bird Index since 1990, underscoring the need for landscape-scale restoration (EEA_FBI_2024). Wild pollinators stand at the centre of the revised EU Pollinators Initiative (“A new deal for pollinators”), which commits all Member States to continent-wide monitoring and trend reversal by 2030 (EU_Pollinators_2023). IAS require surveillance and early-warning systems under Regulation (EU) 1143/2014, which obliges Member States to establish national monitoring and rapid-response mechanisms (EU_IAS_Regulation_2014). Finally, herpetofauna and Mediterranean endemics receive targeted LIFE funding and are focal taxa in Article 17 conservation-status reporting under the Habitats Directive, anchoring them in the EU’s mandatory assessment cycle. Table 3 lists in detail all the species from the above categories that are used for training BioAnalyst.
Table 3: Approximate total occurrences (2000–2020) and GBIF species ID to scientific name mappings for six major species categories
@@
Invasive & alien 8002952 Ambrosia artemisiifolia (common ragweed) 87 k
@@
Mediterranean 2435261 Lynx pardinus (Iberian lynx) 436 k
@@
2434779 Monachus monachus (Med. monk seal) 137 k
@@
Table 4: ERA5 variable names, definitions and units (c3s2017era5; hersbach2020era5)
Variable Units Description
swvl1 $\mathrm{m^{3}\,m^{-3}}$ Volumetric soil water content, layer 1 (0 cm depth)
swvl2 $\mathrm{m^{3}\,m^{-3}}$ Volumetric soil water content, layer 2 (7 cm depth)
@@
smlt m (water eq.) Snow melt accumulated at surface
@@
csfr $\mathrm{kg\,m^{-2}\,s^{-1}}$ Convective snowfall rate
avg_sdswrf $\mathrm{W\,m^{-2}}$ Mean surface downwelling shortwave radiation flux
avg_snswrf $\mathrm{W\,m^{-2}}$ Mean surface net shortwave radiation flux
avg_snlwrf $\mathrm{W\,m^{-2}}$ Mean surface net longwave radiation flux
avg_tprate $\mathrm{m\,s^{-1}}$ Mean total precipitation rate
avg_sdswrfcs $\mathrm{W\,m^{-2}}$ Mean surface downwelling shortwave flux (clear sky)
@@
u $\mathrm{m\,s^{-1}}$ Eastward wind component (pressure levels)
v $\mathrm{m\,s^{-1}}$ Northward wind component (pressure levels)
u10 $\mathrm{m\,s^{-1}}$ 10 m eastward wind component
v10 $\mathrm{m\,s^{-1}}$ 10 m northward wind component
q $\mathrm{kg\,kg^{-1}}$ Specific humidity (pressure levels)
lsm 0/1 Land–sea mask (0 = sea, 1 = land)
avg_slhtf $\mathrm{W\,m^{-2}}$ Mean surface latent heat flux
avg_pevr $\mathrm{kg\,m^{-2}\,s^{-1}}$ Mean potential evaporation rate
@@ B.3 Build Batches
Every input dataset is preprocessed and harmonised based on the following principles:
@@
  • Timestamps are standardised to datetime64 objects with monthly resolution.
@@
To maintain modularity and extensibility, the pipeline separates data loading logic per modality (e.g., _load_era5, _load_csv, _load_species), with all outputs standardised into spatially and temporally aligned xarray.Dataset objects. This design supports the seamless addition of future data types such as Sentinel-2 imagery, elevation, or anthropogenic indicators without altering the core batch assembly logic.
The batching pipeline is optimised to handle large-scale datasets efficiently. ERA5 NetCDF files are processed using xarray with chunking enabled, and CSV files are filtered and indexed spatially using precomputed $0.25^{\circ}$ grid maps. This architecture allows scaling to continental datasets and long temporal ranges without excessive memory usage.
@@
All variables are converted into PyTorch tensors. Scalar geospatial variables (e.g., temperature, NDVI) are stored in tensors of shape $(2,H,W)$, where $H$ and $W$ are the grid height and width and 2 accounts for the consecutive timestamps. Pressure-level variables (e.g., geopotential, wind at altitude) are stored as $(2,C,H,W)$ where $C$ is the number of pressure levels. Species presence tensors are organised as one tensor per species with shape $(2,H,W)$. In cases where pressure-level variables are included, their corresponding pressure levels are saved in the batch metadata.
Each batch includes a metadata dictionary capturing the timestamp window, grid specification (latitude and longitude), list of included species, and pressure levels (if applicable). The final batch is a structured dictionary with modality keys representing the various variable categories (e.g., surface_variables, agriculture_variables, species_variables), each containing variable-specific tensors. The complete batch is serialised and saved to disk in PyTorch’s binary .pt format, enabling efficient loading during model training without repeating and preprocessing operation.
Table 5: Statistics of the constructed data batches used in the pre-training of BioaAnalyst
@@
Grid sampling resolution $0.25^{\circ}$
Global grid size $1440$ (lon) $\times$ $720$ (lat) == $1.036.800$ cells
Europe grid size $280$ (lon) $\times$ $160$ (lat) == $44.800$ cells
Latitude bounds $32^{\circ},\,72^{\circ}]$
Longitude bounds $25^{\circ},\,45^{\circ}]$